OpenStudy (anonymous):

3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

First reminder the student about order of operations. 3-5*(x+2)=3*(x+1)+6 first step multiply the coefficients of the parenthetical expressions first. 3-5x-10=3x+3+6 Next collect similar terms. -5x-7=3x+9 Next move the variables to the same side. +5x=+5x 7=8x+9 now move the nine across 7-9=8x -2=8x -2/8=x -1/4=x

sam (.sam.):

Move all terms not containing x to the right side of the equation. -5(x+2)=3x+6 Multiply -5 by each term inside the parentheses. -5x-10=3x+6 Move terms not containing x to the right side of the equation. -8x-10=6 Move all terms not containg x to the righthand side of the equation. -8x=16 Divide each term in the equation by -8. x=-2

OpenStudy (anonymous):

You distribute the numbers outside the parenthesis first: 3-5x-10=3x+3+6 now you add up like term on each side: -7-5x=3x+9 add five x to each side: -7=8x+9 and subtract 9 from each side: -16=8x now isolate the x by dividing both sides by 8: -2=x ta dah!

OpenStudy (anonymous):

@Mad is kinda similar to mines. but i showed more work :]

OpenStudy (anonymous):

Thank you all so much!

OpenStudy (anonymous):

x=-2